### v1

@@ 2.1 Notation
Mathematical Notation: Vectors are denoted by bold lowercase letters (e.g., $\mathbf{u}$, $\mathbf{v}$), and matrices are denoted by uppercase letters (e.g., $A$, $T$). The horizontal concatenation of matrices is denoted as $[A\mid T]$ or $[A\parallel T]$. The ring of integers is denoted by $\mathbb{Z}$, the ring of integers modulo $q$ by $\mathbb{Z}_{q}$ (or $\mathbb{Z}/q\mathbb{Z}$), a general field by $\mathbb{F}$, and the finite field with $q$ elements (where $q$ is a prime or a prime power) by $\mathbb{F}_{q}$. The $n$-dimensional Euclidean space is denoted by $\mathbb{R}^{n}$.
@@
[…]
Asymptotic Notation: We employ the notations $\mathcal{O}$, $\widetilde{\mathcal{O}}$, $\Omega$, $\Theta$, and $\omega$, along with $\text{poly}(n)$, to describe the asymptotic growth relations of functions. For two functions $f(n)$ and $g(n)$, $\mathcal{O}$ denotes the asymptotic upper bound of $f(n)$, meaning there exist positive constants $c$ and $N$ such that for all $n\geq N$, $0\leq f(n)\leq c\cdot g(n)$. The notation $\widetilde{\mathcal{O}}$ ignores logarithmic factors, indicating that $f(n)=\mathcal{O}(g(n)\cdot\log^{k}n)$ holds for some $k$. The symbol $\Omega$ represents the asymptotic lower bound, implying the existence of $c>0$ and $N$ such that for $n\geq N$, $0\leq c\cdot g(n)\leq f(n)$. The notation $\Theta$ signifies a tight bound, meaning both $\mathcal{O}$ and $\Omega$ are satisfied simultaneously. The notation $\omega$ is a stricter form of $\Omega$, indicating that $\lim_{n\to\infty}\frac{f(n)}{g(n)}=\infty$. Lastly, $\text{poly}(n)$ denotes polynomial growth, i.e., $f(n)=\mathcal{O}(n^{k})$ for some $k$. A function $f(n)$ is called negligible, denoted as $\mathrm{negl}(n)$, if for every constant $c>0$, $f(n)$ is $O(n^{-c})$. Moreover, an event is said to occur with overwhelming probability if its probability is $1-\mathrm{negl}(n)$.
Probabilistic Notation and Algorithms: For any given distribution $\mathcal{D}$, the expression $x\sim\mathcal{D}$ signifies that the variable $x$ follows the distribution $\mathcal{D}$, while $x\leftarrow\mathcal{D}$ denotes the process of sampling a random value from $\mathcal{D}$. When considering a finite set $\mathcal{X}$, the notation $x\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathcal{X}$ indicates that $x$ is chosen uniformly at random from $\mathcal{X}$. Regarding probabilistic polynomial-time (PPT) algorithms $\mathsf{Alg}$, the notation $y\leftarrow\mathsf{Alg}(x)$ represents the procedure of generating an output $y$ by executing the algorithm on input $x$.
@@ 2.2 Statistical Distance, Entropy and Decomposition Algorithm
is a negligible function in the security parameter $\lambda$.
@@
$\mathcal{X}$ =$=\big{\{}x\in\mathbb{F}_{2}^{2k}\mid\|x\|=k-1\big{\}}\sqcup\{0\},$
$\mathcal{Y}$ =$=\big{\{}y\in\mathbb{F}_{2}^{2k}\mid\|y\|=k\big{\}}\sqcup\{0\},$
$\mathcal{Z}$ =$=\big{\{}z\in\mathbb{F}_{2}^{2k}\mid\|z\|=k+1\big{\}}\sqcup\{0\}.$
@@ 2.3 Background on Lattices and Hard Problems
(1)$\Lambda_{q}^{\bot}=\{\mathbf{x}\in\mathbb{Z}^{m}|A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q)\}$;
(2)$\Lambda_{q}^{\mathbf{u}}=\{\mathbf{y}\in\mathbb{Z}^{m}|A\cdot\mathbf{y}\equiv\mathbf{u}(\mathrm{mod}q)\}$.
@@ 2.4 Gaussian Distribution and Its Related Algorithms
[…]
@@
    There is a probabilistic polynomial-time algorithm SampleGaussian$(A,T,s,\mathbf{c})$ that outputs $\mathbf{x}\in\Lambda_{q}^{\bot}(A)$ drawn from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\bot}(A),s,\mathbf{c}}$;
@@
    There is a probabilistic polynomial-time algorithm SamplePre$(A,T,\mathbf{u},s)$ that outputs $\mathbf{x}\in\Lambda_{q}^{\mathbf{u}}(A)$ sampled from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\mathbf{u}}(A),s}$, whenever $\Lambda_{q}^{\mathbf{u}}(A)$ is not empty.
@@
([38]) Let $n$ and $h$ be positive integers, and let $q$ be a prime satisfying $n\geq 2h\log q$. Then, for all matrices $A\in\mathbb{Z}_{q}^{h\times n}$ except a negligible fraction of $2q^{-h}$, and for any $s\geq\omega(\sqrt{\log n})$, the distribution of $\mathbf{\alpha}=A\cdot\mathbf{x}\pmod{q}$ is statistically indistinguishable from uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}$ is sampled as $\mathbf{x}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textsf{SampleDom}(1^{n},s)$.
@@
([38]) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$
@@ 2.5 Statistical Properties Related to Discrete Gaussian Distributions
3 Definition and Security Model of Linearly Homomorphic Ring Signature
3.1 Definition of Linearly Homomorphic Ring Signature
@@
(LHRS scheme) The linearly homomorphic ring signature is composed of a set of probabilistic polynomial-time (PPT) algorithms $\mathcal{LHRS}$ =(KeyGen,HR-Sign,Combine,HR-Verify), which are defined as follows:
@@ 1. $\bullet$
    KeyGen ($1^{\lambda}$, pp): Given a security parameter $n$ and public parameters pp, this algorithm outputs a key pair $(pk,sk)$. Here, the public parameters pp determine the tag space $\mathcal{T}$ , the message space $\mathcal{M}$, the signature space $\Sigma$, and and the maximum number of homomorphic operations $k$.
@@ 2. $\bullet$
    HR-Sign$(pk_{i},sk_{i},R,\tau,\mathbf{m})$: Given the key pair $(pk_{i},sk_{i})$ of user $i$, the ring’s public key set $R$ (where $pk_{i}\in R$), the dataset label $\tau\in\mathcal{T}$, and the message $\mathbf{m}\in\mathcal{M}$, this algorithm outputs a ring signature $\sigma$ on message $\mathbf{m}$ under label $\tau$, signed using $sk_{i}$.
@@ 3. $\bullet$
    Combine$(R,\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{l})$: Given the ring $R$, the tag $\tau$, and a set of tuples $\{(c_{j},\sigma_{j})\}_{j=1}^{l})$ (where $l\leq k$), this algorithm outputs a ring signature $\sigma$ on the message $\sum_{i=j}^{l}c_{j}\mathbf{m}_{j}$ under the label tag $\tau$ for the ring $R$
@@ 4. $\bullet$
    HR-Verify$(R,\tau,\mathbf{m},\sigma)$: Given the ring $R$, the tag $\tau$, the message $\mathbf{m}$, and the ring signature $\sigma$, the algorithm outputs 0 (reject) or 1 (accept).
@@
(1) For all valid tags $\tau\in\mathcal{T}$ and messages $\mathbf{m}\in\mathcal{M}$, if a ring signature $\sigma$ is generated as $\sigma\leftarrow\textsf{HR-Sign}(pk_{i},sk_{i},R,\tau,\mathbf{m})$, then the verification algorithm must satisfy
[…]
(2) If $\sigma_{j}\leftarrow\textsf{HR-Sign}(pk_{i},sk_{i},R,\tau,\mathbf{m}_{j})$, then
[…]
3.2 Security Model of Linearly Homomorphic Ring Signature
@@ 1. $\bullet$
    Setup: The challenger $\mathcal{B}_{1}$ runs the $\mathsf{Setup}(1^{\lambda},\mathsf{pp})$ algorithm $l$ times to generate key pairs $(pk_{1},sk_{1}),\dots,(pk_{l},sk_{l})$, where $l$ is a game parameter, and sends the public key set $\{pk_{s}\}_{s=1}^{l}$ to the adversary $\mathcal{A}_{1}$.
@@ 2. $\bullet$
    Queries: After receiving $\{(pk_{s},sk_{s})\}_{s=1}^{l}$, the adversary $\mathcal{A}_{1}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{1}$ selects a user $s\in[l]=\{1,2,...,l\}$ and sends $s$ to the challenger $\mathcal{B}_{1}$, who responds by returning $sk_{s}$ to $\mathcal{A}_{1}$.
    -Sign query: When $\mathcal{A}_{1}$ queries a ring signature on basis vectors $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k}$ of subspace $V_{si}$ for user $s$ with $pk_{s}\in R$, the challenger $\mathcal{B}_{1}$ first generates a uniformly random tag $\tau_{si}\xleftarrow{\$}\mathcal{T}$ for $V_{si}$ and responds with tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for $i=1,...,q_{s}$ and $j=1,...,k$.
@@ 3. $\bullet$
    Challenge: Finally, $\mathcal{A}_{1}$ outputs a challenge tuple $(s_{0},s_{1},R^{*},\tau^{*},m^{*})$ to $\mathcal{B}_{1}$, where $s_{0}$ and $s_{1}$ are indices satisfying $pk_{s_{0}}\in R^{*}$ and $pk_{s_{1}}\in R^{*}$, requiring $\mathcal{B}_{1}$ to sign $(\tau^{*},m^{*})$ under ring $R^{*}$. $\mathcal{B}_{1}$ then selects a random bit $b\xleftarrow{\$}\{0,1\}$, computes $\sigma_{b}^{*}\leftarrow\textsf{HR-Sign}(pk_{s_{b}},sk_{s_{b}},R^{*},\tau^{*},m^{*})$, and returns $\sigma_{b}^{*}$ to $\mathcal{A}_{1}$. The adversary outputs a guess $b^{\prime}\in\{0,1\}$ and wins the game if $b^{\prime}=b$.
@@ 1. $\bullet$
    Setup: The challenger $\mathcal{B}_{2}$ runs the $\mathsf{Setup}(1^{\lambda},\mathsf{pp})$ algorithm $l$ times to generate key pairs $(pk_{1},sk_{1}),\dots,(pk_{l},sk_{l})$, where $l$ is a game parameter. Let $L=\{pk_{s}\}_{s=1}^{l}$ and sends the public key set $L$ to the adversary $\mathcal{A}_{2}$
@@ 2. $\bullet$
    Queries: After receiving $L=\{pk_{s}\}_{s=1}^{l}$, the adversary $\mathcal{A}_{2}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{2}$ selects a user $s\in[l]=\{1,2,...,l\}$ and sends $s$ to the challenger $\mathcal{B}_{2}$, who responds by returning $sk_{s}$ to $\mathcal{A}_{2}$.
    -Signing query: When $\mathcal{A}_{2}$ queries a ring signature on basis vectors $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k}$ of subspace $V_{si}$ for user $s$ with $pk_{s}\in R$, the challenger $\mathcal{B}_{2}$ first generates a uniformly random tag $\tau_{si}\xleftarrow{\$}\mathcal{T}$ for $V_{si}$ and responds with tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for $i=1,...,q_{s}$ and $j=1,...,k$.

### v3

@@ 2.1 Notation
Mathematical Notation: Vectors are denoted by bold lowercase letters (e.g., $\mathbf{u}$, $\mathbf{v}$), and matrices are denoted by uppercase letters (e.g., $A$, $T$). The horizontal concatenation of matrices is denoted as $A\mid T]$ or $A\parallel T]$. The ring of integers is denoted by $\mathbb{Z}$, the ring of integers modulo $q$ by $\mathbb{Z}_{q}$ (or $\mathbb{Z}/q\mathbb{Z}$), a general field by $\mathbb{F}$, and the finite field with $q$ elements (where $q$ is a prime or a prime power) by $\mathbb{F}_{q}$. The $n$-dimensional Euclidean space is denoted by $\mathbb{R}^{n}$.
@@
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product in Euclidean space.
Asymptotic Notation: For asymptotic analysis, we employ the following notation:
  • $f(n)=\mathcal{O}(g(n))$ if $\exists c>0,N\in\mathbb{N}$ such that $\forall n>N$, $f(n)\leq cg(n)$
  • $f(n)=\widetilde{\mathcal{O}}(g(n))$ when $f(n)=\mathcal{O}(g(n)\log^{c^{\prime}}n)$ for some $c^{\prime}>0$
  • $f(n)=\omega(g(n))$ if $\forall c>0$, $\exists N$ such that $g(n)\leq cf(n)$ whenever $n>N$
Let $n$ be the security parameter. A function $f(n)$ is said to be polynomial (written $f(n)=\mathrm{poly}(n)$) if there exists $c>0$ such that $f(n)=\mathcal{O}(n^{c})$; conversely, it is negligible (denoted $\mathrm{negl}(n)$) if for every $c>0$ we have $f(n)=\mathcal{O}(n^{-c})$. An event is said to occur with overwhelming probability when its probability satisfies $\Pr[\text{event}]\geq 1-\mathrm{negl}(n)$.
Probabilistic Notation and Algorithms: For any given distribution $\mathcal{D}$, the expression $x\sim\mathcal{D}$ signifies that the variable $x$ follows the distribution $\mathcal{D}$, while $x\leftarrow\mathcal{D}$ denotes the process of sampling a random value from $\mathcal{D}$. When considering a finite set $\mathcal{X}$, the notation $x\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{X}$ indicates that $x$ is chosen uniformly at random from $\mathcal{X}$. Regarding probabilistic polynomial-time (PPT) algorithms $\mathsf{Alg}$, the notation $y\leftarrow\mathsf{Alg}(x)$ represents the procedure of generating an output $y$ by executing the algorithm on input $x$.
@@ 2.2 Statistical Distance, Entropy and Decomposition Algorithm
is a negligible function in the security parameter $n$.
@@
$\mathcal{X}$ $=\big\{x\in\mathbb{F}_{2}^{2k}\mid\|x\|=k-1\big\}\sqcup\{0\},$
$\mathcal{Y}$ $=\big\{y\in\mathbb{F}_{2}^{2k}\mid\|y\|=k\big\}\sqcup\{0\},$
$\mathcal{Z}$ $=\big\{z\in\mathbb{F}_{2}^{2k}\mid\|z\|=k+1\big\}\sqcup\{0\}.$
@@ 2.3 Background on Lattices and Hard Problems
[…]
@@ 2.4 Gaussian Distribution and Its Related Algorithms
([39]) Let $q\geq 3$ be odd, $n$ be a positive integer, and let $m:=\lceil 6n\log q\rceil.$ There is a probabilistic polynomial-time algorithm TrapGen(q,n,m)(q,n,m) that outputs a pair $(A,T)$ such that $A$ is statistically close to a uniform rank $n$ matrix in $\mathbb{Z}_{q}^{n\times m}$, and $T\in\mathbb{Z}^{n\times n}$ is a basis for $\Lambda_{q}^{\bot}(A)$ satisfying
@@
[…]
@@
[…]
@@
([38]) Let $n$ and $h$ be positive integers, and let $q$ be a prime satisfying $n\geq 2h\log q$. Then, for all matrices $A\in\mathbb{Z}_{q}^{h\times n}$ except a negligible fraction of $2q^{-h}$, and for any $s\geq\omega(\sqrt{\log n})$, the distribution of $\mathbf{\alpha}=A\cdot\mathbf{x}\pmod{q}$ is statistically indistinguishable from uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}$ is sampled as $\mathbf{x}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textsf{SampleDom}(1^{n},s)$.
@@
([38], Lemma 5.3) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$ .
@@ 2.5 Statistical Properties Related to Discrete Gaussian Distributions
3 Definition and Security Model of LHRS
3.1 Definition of LHRS
@@
(LHRS scheme) The linearly homomorphic ring signature is composed of a set of probabilistic polynomial-time (PPT) algorithms $\mathcal{LHRS}$ =(Setup,HR-Sign,Combine,HR-Verify), which are defined as follows:
@@ 1. $\bullet$
    Setup ($1^{n}$, pp): Given a security parameter $n$ and public parameters pp, this algorithm outputs a key pair $(pk,sk)$. Here, the public parameters pp determine the tag space $\mathcal{T}$ , the message space $\mathcal{M}$, the signature space $\Sigma$, and and the maximum number of homomorphic operations $k_{0}$.
@@ 2. $\bullet$
[…]
@@ 3. $\bullet$
[…]
@@ 4. $\bullet$
[…]
@@
(1) For all valid tags $\tau\in\mathcal{T}$ and messages $\mathbf{m}\in\mathcal{M}$, if a ring signature $\sigma$ is generated as $\sigma\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R,\tau,\mathbf{m})$, then the verification algorithm must satisfy
$1\leftarrow\textsf{HR-Verify}(R,\tau,\mathbf{m},\sigma).$
(2) If $\sigma_{j}\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R,\tau,\mathbf{m}_{j})$, then
$1\leftarrow\textsf{HR-Verify}(R,\tau,\sum_{j=1}^{\ell}c_{j}\mathbf{m}_{j},\textsf{Combine}(R,\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{\ell})).$
3.2 Security Model of LHRS
@@ 1. $\bullet$
    Setup: The challenger $\mathcal{B}_{1}$ runs the $\mathsf{Setup}(1^{n},\mathsf{pp})$ algorithm $\ell$ times to generate key pairs $(\textsf{pk}_{1},\textsf{sk}_{1}),\dots,(\textsf{pk}_{\ell},\textsf{sk}_{\ell})$, where $\ell$ is a game parameter, and sends the public key set $\{\textsf{pk}_{s}\}_{s=1}^{\ell}$ to the adversary $\mathcal{A}_{1}$.
@@ 2. $\bullet$
    Queries: After receiving $\{(\textsf{pk}_{s},\textsf{sk}_{s})\}_{s=1}^{\ell}$, the adversary $\mathcal{A}_{1}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{1}$ selects a user $s\in[\ell]=\{1,2,$$\ell\}$ and sends $s$ to the challenger $\mathcal{B}_{1}$, who responds by returning $\textsf{sk}_{s}$ to $\mathcal{A}_{1}$.
    -Sign query: After selecting a user $s$, adversary $\mathcal{A}_{2}$ adaptively selects a series of message subspaces $V_{si}=\text{span}\{\mathbf{m}_{si}^{(1)},...,\mathbf{m}_{si}^{(k_{0})}\}$ and queries them to the challenger $\mathcal{B}_{2}$, for $i=1,\dots,q_{s}$, where $q_{s}$ is the maximum number of allowed queries for user $s$. For each queried subspace $V_{si}$, the challenger $\mathcal{B}_{2}$ generates a uniformly random tag $\tau_{si}\xleftarrow{\mathdollar}\mathcal{T}$. Then, $\mathcal{B}_{2}$ computes a signature $\sigma_{si}^{(j)}$ for each message $\mathbf{m}_{si}^{(j)}$ within the subspace, in relation to the tag $\tau_{si}$. Finally, $\mathcal{B}_{2}$ responds with the tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for all $i=1,...,q_{s}$ and $j=1,\dots,k_{0}$.
@@ 3. $\bullet$
    Challenge: Finally, $\mathcal{A}_{1}$ outputs a challenge tuple $(s_{0},s_{1},R^{*},\tau^{*},m^{*})$ to $\mathcal{B}_{1}$, where $s_{0}$ and $s_{1}$ are indices satisfying $\textsf{pk}_{s_{0}}\in R^{*}$ and $\textsf{pk}_{s_{1}}\in R^{*}$, and neither $s_{0}$ nor $s_{1}$ has been corrupted in any prior corruption query. $\mathcal{B}_{1}$ is then required to generate a signature on $(\tau^{*},m^{*})$ under ring $R^{*}$. To do so, $\mathcal{B}_{1}$ selects a random bit $b\xleftarrow{\mathdollar}\{0,1\}$, computes $\sigma_{b}^{*}\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s_{b}},\textsf{sk}_{s_{b}},R^{*},\tau^{*},m^{*})$, and returns $\sigma_{b}^{*}$ to $\mathcal{A}_{1}$. The adversary outputs a guess $b^{\prime}\in\{0,1\}$ and wins the game if $b^{\prime}=b$.
@@ 1. $\bullet$
    Setup: The challenger $\mathcal{B}_{2}$ runs the $\mathsf{Setup}(1^{n})$ algorithm $\ell$ times to generate key pairs $(\textsf{pk}_{1},\textsf{sk}_{1}),\dots,(\textsf{pk}_{\ell},\textsf{sk}_{\ell})$, where $\ell$ is a game parameter. Let $L=\{\textsf{pk}_{s}\}_{s=1}^{\ell}$ and sends the public key set $L$ to the adversary $\mathcal{A}_{2}$
@@ 2. $\bullet$
    Queries: After receiving $L=\{\textsf{pk}_{s}\}_{s=1}^{\ell}$, the adversary $\mathcal{A}_{2}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{2}$ selects a user $s\in[\ell]=\{1,2,$$\ell\}$ and sends $s$ to the challenger $\mathcal{B}_{2}$, who responds by returning $\textsf{sk}_{s}$ to $\mathcal{A}_{2}$.
    -Signing query: After selecting a user $s$, adversary $\mathcal{A}_{2}$ adaptively selects a series of message subspaces $V_{si}=\text{span}\{\mathbf{m}_{si}^{(1)},...,\mathbf{m}_{si}^{(k_{0})}\}$ and queries them to the challenger $\mathcal{B}_{2}$, for $i=1,\dots,q_{s}$, where $q_{s}$ is the maximum number of allowed queries for user $s$. For each queried subspace $V_{si}$, the challenger $\mathcal{B}_{2}$ generates a uniformly random tag $\tau_{si}\xleftarrow{\mathdollar}\mathcal{T}$. Then, $\mathcal{B}_{2}$ computes a signature $\sigma_{si}^{(j)}$ for each message $\mathbf{m}_{si}^{(j)}$ within the subspace, in relation to the tag $\tau_{si}$. Finally, $\mathcal{B}_{2}$ responds with the tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for all $i=1,...,q_{s}$ and $j=1,\dots,k_{0}$.